# The geodesic cover problem for butterfly networks

### v2

@@ 3.2 An upper bound for the geodesic cover number of butterfly networks
As in Case $1$, we we can construct $34\ell+$ 3 geodesics to cover all the vertices of $U^{r}$, $U^{b}$, $W^{r}$, and $W^{b}$. We have thus constructed $34\ell+$ 3 $=\lceil\frac{2^{r-1}}{3}\rceil$ geodesics covering all the vertices of $U^{r}\cup U^{b}\cup W^{r}\cup W^{b}=A^{\prime}\cup B^{\prime}\cup C^{\prime}\cup D^{\prime}$.
@@ 4 The edge geodesic cover problem
Lemma 4.1.
If $r\geq 3$, then $E({\rm BF}(r))$ can be partitioned by a set $S(r)$ of edge-disjoint isometric cycles of length $4r$, where $|S(r)|=2^{r-1}$ and each isometric cycle of $S(r)$ has two vertices at level 00.
Proof.
The proof is by induction on the dimension $r$ of ${\rm BF}(r)$. The base case is $r=3$ and is ellaborated in Fig. 7, where $E({\rm BF}(3))$ is partitioned by a set $S(3)$ of edge-disjoint isometric cycles of length $4\cdot 3$ and each isometric cycle of $S(3)$ has two vertices at level 00.
@@
Assume now that the edge set of ${\rm BF}(k-1)$ can be partitioned by a set $S(k-1)$ of edge-disjoint isometric cycles of length $4(k-1)$, where $|S|=2^{k-2}$ and each isometric cycle $C$ of $S(k-1)$ has two vertices $u$ and $w$ at level 00. A cycle $C$ in $S(k-1)$ is represented by $u-P-w-Q-u$ where $u$ and $w$ are the two vertices of $C$ at level 00, $P$ is the path segment of length $22k-$ 2 in $C$ between $u$ and $w$ and $Q$ is the path segment of length $22k-$ 2 in $C$ between $u$ and $w$. Refer to Fig. 8.
Figure 8: (a) An isometric cycle $C=u-P-w-Q-u$ in ${\rm BF}(k-1)$. (b) Two copies of $C$ are $C^{\prime}$ and $C^{\prime\prime}$. The two isometric cycles $C^{\prime}$ and $C^{\prime\prime}$ are connected by vertices $a$, $b$, $x$ and $y$ which are at level 00. (c) The two isometric cycles $C^{\prime}$ and $C^{\prime\prime}$ generate two different isometric cycles $C_{1}$ and $C_{2}$ (marked in different color) in ${\rm BF}(k)$.
Recall that ${\rm BF}(r)$ has two copies of ${\rm BF}(k-1)$, ${\rm BF}^{\prime}(k-1)$ and ${\rm BF}^{\prime\prime}(k-1)$, where $V({\rm BF}^{\prime}(k-1))=\{v^{\prime}:\ v\in{\rm BF}(k-1)\}$ and $V({\rm BF}^{\prime\prime}(k-1))=\{v^{\prime\prime}:\ v\in{\rm BF}(k-1)\}$. Also, there are two copies of $S(k-1)$ in ${\rm BF}(k)$, $S^{\prime}(k-1)$ and $S^{\prime\prime}(k-1)$, where $S^{\prime}(k-1)=\{C^{\prime}=u^{\prime}-P^{\prime}-w^{\prime}-Q^{\prime}-u^{\prime}:\ C=u-P-w-Q-u\in S(k-1)\}$ and $S^{\prime\prime}(k-1)=\{C^{\prime\prime}=u^{\prime\prime}-P^{\prime\prime}-w^{\prime\prime}-Q^{\prime\prime}-u^{\prime\prime}:\ C=u-P-w-Q-u\in S(k-1)\}$. The length of cycles $C\in S(k-1)$, $C^{\prime}\in S^{\prime}(k-1)$, and $C^{\prime\prime}\in S^{\prime\prime}(k-1)$ is $4(k-1)$. In ${\rm BF}(r)$, there are two vertices $a$ and $b$ at level 00 which are adjacent to $u^{\prime}$ of $C^{\prime}$ and $u^{\prime\prime}$ of $C^{\prime\prime}$. In the same way, there are two vertices $x$ and $y$ at level 00 which are adjacent to $w^{\prime}$ of $C^{\prime}$ and $w^{\prime\prime}$ of $C^{\prime\prime}$. Refer to Fig. 8. Now define the cycles $C_{1}=a-u^{\prime}-P^{\prime}-w^{\prime}-x-w^{\prime\prime}-P^{\prime\prime}-u^{\prime\prime}-a$ and $C_{2}=b-u^{\prime}-Q^{\prime}-w^{\prime}-y-w^{\prime\prime}-Q^{\prime\prime}-u^{\prime\prime}-b$. Since, $P^{\prime}$, $P^{\prime\prime}$, $Q^{\prime}$, and $Q^{\prime\prime}$ each are of length $22k-$ 2, it is easy to observe that the length of $C_{1}$ and $C_{2}$ is $4k$. Let us define $S(k)=\{C_{1},C_{2}:C\in S(k-1)\}$. Each cycle $C_{1}$ and $C_{2}$ have two vertices at level 00. It is easy to observe that the cycles in $S(k)$ are isometric and they are mutually edge-disjoint. Moreover, the cardinality of $S(k)$ is $2^{k-1}$. Thus, the edge set of ${\rm BF}(k)$ is partitioned by the edge-disjoint isometric cycles of $S(k)$ such that $|S(k)|=2^{k-1}$. ∎
Theorem 4.2.
@@
Proof.
In order to derive the claimed lower bound for ${\rm gpart_{e}}({\rm BF}(r))$, let us consider all $(2,4)$-edges of ${\rm BF}(r)$. As already mentioned, there are $2^{r+2}$ $(2,4)$-edges in ${\rm BF}(r)$. Since a geodesic can cover a maximum of four $(2,4)$-edges of ${\rm BF}(r)$, ${\rm gcover_{e}}({\rm BF}(r))\geq 2^{r+2}/4=2^{r}$. Thus, ${\rm gpart_{e}}({\rm BF}(r))\geq{\rm gcover_{e}}({\rm BF}(r))\geq 2^{r}$.
To prove ${\rm gpart_{e}}({\rm BF}(r))\leq 2^{r}$, it is enough to construct an edge geodesic partition of cardinality $2^{r}$ for ${\rm BF}(r)$. By Lemma 4.1, the edge set of ${\rm BF}(r)$ can be partitioned by a set $S$ of edge-disjoint isometric cycles of length $4r$, where $|S|=2^{r-1}$. In other words, the edge set of ${\rm BF}(r)$ can be partitioned by a set $R$ of edge-disjoint diametrals of length $2r$ such that $|R|=2^{r}$. Thus, ${\rm gpart_{e}}({\rm BF}(r))\leq 2^{r}$. ∎

### v3

@@ 3.2 An upper bound for the geodesic cover number of butterfly networks
As in Case $1$, we we can construct $4\ell+3$ geodesics to cover all the vertices of $U^{r}$, $U^{b}$, $W^{r}$, and $W^{b}$. We have thus constructed $4\ell+3$ $\lceil\frac{2^{r-1}}{3}\rceil$ geodesics covering all the vertices of $U^{r}\cup U^{b}\cup W^{r}\cup W^{b}=A^{\prime}\cup B^{\prime}\cup C^{\prime}\cup D^{\prime}$.
@@ 4 The edge geodesic cover problem
Lemma 4.1
If $r\geq 3$, then $E({\rm BF}(r))$ can be partitioned by a set $S(r)$ of edge-disjoint isometric cycles of length $4r$, where $|S(r)|=2^{r-1}$ and each isometric cycle of $S(r)$ has two vertices at level 0.
Proof 4.2
The proof is by induction on the dimension $r$ of ${\rm BF}(r)$. The base case is $r=3$ and is ellaborated in Fig. 7, where $E({\rm BF}(3))$ is partitioned by a set $S(3)$ of edge-disjoint isometric cycles of length $4\cdot 3$ and each isometric cycle of $S(3)$ has two vertices at level 0.
@@
Assume now that the edge set of ${\rm BF}(k-1)$ can be partitioned by a set S(k1)S(k-1) of edge-disjoint isometric cycles of length 4(k1)4(k-1), where $|S|=2^{k-2}$ and each isometric cycle $C$ of S(k1)S(k-1) has two vertices $u$ and $w$ at level 0. A cycle $C$ in S(k1)S(k-1) is represented by uPwQuu-P-w-Q-u where $u$ and $w$ are the two vertices of $C$ at level 0, $P$ is the path segment of length 2k22k-2 in $C$ between $u$ and $w$ and $Q$ is the path segment of length 2k22k-2 in $C$ between $u$ and $w$. Refer to Fig. 8.
Figure 8: (a) An isometric cycle C=uPwQuC=u-P-w-Q-u in ${\rm BF}(k-1)$. (b) Two copies of $C$ are $C^{\prime}$ and $C^{\prime\prime}$. The two isometric cycles $C^{\prime}$ and $C^{\prime\prime}$ are connected by vertices $a$, $b$, $x$ and $y$ which are at level 0. (c) The two isometric cycles $C^{\prime}$ and $C^{\prime\prime}$ generate two different isometric cycles $C_{1}$ and $C_{2}$ (marked in different color) in ${\rm BF}(k)$.
Recall that ${\rm BF}(r)$ has two copies of ${\rm BF}(k-1)$, ${\rm BF}^{\prime}(k-1)$ and ${\rm BF}^{\prime\prime}(k-1)$, where $V({\rm BF}^{\prime}(k-1))=\{v^{\prime}:\ v\in{\rm BF}(k-1)\}$ and $V({\rm BF}^{\prime\prime}(k-1))=\{v^{\prime\prime}:\ v\in{\rm BF}(k-1)\}$. Also, there are two copies of S(k1)S(k-1) in ${\rm BF}(k)$, $S^{\prime}(k-1)$ and $S^{\prime\prime}(k-1)$, where
$S^{\prime}(k-1)=\{C^{\prime}=u^{\prime}-P^{\prime}-w^{\prime}-Q^{\prime}-u^{\prime}:\ C=u-P-w-Q-u\in S(k-1)\}$
and
$S^{\prime\prime}(k-1)=\{C^{\prime\prime}=u^{\prime\prime}-P^{\prime\prime}-w^{\prime\prime}-Q^{\prime\prime}-u^{\prime\prime}:\ C=u-P-w-Q-u\in S(k-1)\}.$
The length of cycles $C\in S(k-1)$, $C^{\prime}\in S^{\prime}(k-1)$, and $C^{\prime\prime}\in S^{\prime\prime}(k-1)$ is 4(k1)4(k-1). In ${\rm BF}(r)$, there are two vertices $a$ and $b$ at level 0 which are adjacent to $u^{\prime}$ of $C^{\prime}$ and $u^{\prime\prime}$ of $C^{\prime\prime}$. In the same way, there are two vertices $x$ and $y$ at level 0 which are adjacent to $w^{\prime}$ of $C^{\prime}$ and $w^{\prime\prime}$ of $C^{\prime\prime}$. Refer to Fig. 8. Now define the cycles
$C_{1}=a-u^{\prime}-P^{\prime}-w^{\prime}-x-w^{\prime\prime}-P^{\prime\prime}-u^{\prime\prime}-a$
and
$C_{2}=b-u^{\prime}-Q^{\prime}-w^{\prime}-y-w^{\prime\prime}-Q^{\prime\prime}-u^{\prime\prime}-b.$
Since, $^{\prime}, P′′P^{\prime\prime}$, $Q^{\prime}$, and $Q^{\prime\prime}$ each are of length 2k22k-2, it is easy to observe that the length of $C_{1}$ and $C_{2}$ is $4k$. Let us define $S(k)=\{C_{1},C_{2}:C\in S(k-1)\}$. Each cycle $C_{1}$ and $C_{2}$ have two vertices at level 0. It is easy to observe that the cycles in $S(k)$ are isometric and they are mutually edge-disjoint. Moreover, the cardinality of $S(k)$ is $2^{k-1}$. Thus, the edge set of ${\rm BF}(k)$ is partitioned by the edge-disjoint isometric cycles of $S(k)$ such that $|S(k)|=2^{k-1}$.
Theorem 4.3
@@
Proof 4.4
In order to derive the claimed lower bound for ${\rm gpart_{e}}({\rm BF}(r))$, let us consider all $(2,4)$-edges of ${\rm BF}(r)$. As already mentioned, there are $2^{r+2}$ $(2,4)$-edges in ${\rm BF}(r)$. Since a geodesic can cover a maximum of four $(2,4)$-edges of ${\rm BF}(r)$, ${\rm gcover_{e}}({\rm BF}(r))\geq 2^{r+2}/4=2^{r}$. Thus,
${\rm gpart_{e}}({\rm BF}(r))\geq{\rm gcover_{e}}({\rm BF}(r))\geq 2^{r}.$
To prove ${\rm gpart_{e}}({\rm BF}(r))\leq 2^{r}$, it is enough to construct an edge geodesic partition of cardinality $2^{r}$ for ${\rm BF}(r)$. By Lemma 4.1, the edge set of ${\rm BF}(r)$ can be partitioned by a set $S$ of edge-disjoint isometric cycles of length $4r$, where $|S|=2^{r-1}$. In other words, the edge set of ${\rm BF}(r)$ can be partitioned by a set $R$ of edge-disjoint diametrals of length $2r$ such that $|R|=2^{r}$. Thus, ${\rm gpart_{e}}({\rm BF}(r))\leq 2^{r}$.